OpenStudy (anonymous):

Find all of the zeros of f(x) if f(x) = (x2 – 4)3. Identify any multiple zeros. A. {4 (triple zero), –4 (triple zero)} B. {4 (triple zero)} C. {2, –2} D. {2 (triple zero), –2 (triple zero)}

OpenStudy (anonymous):

I am confused what this question is trying to ask

OpenStudy (bahrom7893):

this one's pretty easy: \[(x^2-4)^3=((x-2)(x+2))^3=(x-2)(x+2)(x-2)(x+2)(x-2)(x+2)\]

OpenStudy (anonymous):

x^6 - 64

OpenStudy (anonymous):

But what do they mean by multiple zero

OpenStudy (anonymous):

Wouldn't the answer be D

OpenStudy (bahrom7893):

Multiple 0 basically means how many times that same answer let's you get 0 (i know that kinda doesn't make sense), but basically in our case we'd have a triple zero in the case of 2 and in the case of -2, so yea the answer would be D (refer to my expanded equation)

OpenStudy (anonymous):

Ok thanks